OpenStudy (lilai3):

puhleeze help!!!.................................................sigh....0.12 ft cubed is how many centimeters cubed?

OpenStudy (anonymous):

\[\large 0.12 \cancel{ft^3} \times \frac{12^3 \cancel{in^3}}{\cancel{ft^3}} \times \frac{2.54^3 cm^3}{\cancel{in^3}}\]

OpenStudy (lilai3):

is the answer 3403.715?

OpenStudy (anonymous):

No, but that's close. You might have rounded off something in the middle of your calculation. Don't round any numbers until the final answer.

OpenStudy (lilai3):

you mean that is not the final answer? and what answer did you get? i didn't round it off. it came exactly from my scientific calculator (Texas instruments) [[california usa]]

OpenStudy (anonymous):

\(\large 0.12 \times 12^3 \times 2.54^3 \approx 3398\) which should be rounded to 3400 cm^3

OpenStudy (lilai3):

why do you have to multiply it by 12 to the third power? where did twelve come from? i understand the 2.54 cubed part though cuz there is 2.54 centimeters in one inch and..yeah

OpenStudy (anonymous):

12 inches per foot. 2.54 cm per inch. You have to cube them because you have cubic feet.

OpenStudy (lilai3):

oh i get it now. so is the final answer 3998?

OpenStudy (lilai3):

uhm is that it? cuz i gtg uhm thanks anyway @CliffSedge i got to go bye

OpenStudy (anonymous):

Yeah, that's all you need to know, but the final answer should really be 3400 cm^3 since your given information only had 2 significant figures.

OpenStudy (lilai3):

but on the front you said that ^^^ you should not round the number off until you get the final answer and you got 3998 and then now you said that it should be really rounded off to 63400 what do you mean